Hi! I need help with two problems. (^ means exponent follows)
2e^2x = 5^x+1
and...
log(x+1)-log(x)=2
thanks!

Answer by Nate(3500)

2e^(2x) = 5^(x+1)
e^(2x) = 5^(x+1)/2
2x = (x+1)ln(5)-ln(2)
2x = 1.6x + 1.6 - 0.69
0.4x = 0.91
x=2.275
The number is about 2.275
.............................................................................
log(x+1)-log(x)=2
log((x+1)/x))=2
(x+1)/(x)=100
x+1=100x
1=99x
x=(1/99)